OpenStudy (anonymous):

A jar contains 5 blue, 4 red, 12 white, and 3 green marbles. What is the theoretical probability that a marble drawn at random is NOT white? Express your answer in the form of a percent.

OpenStudy (shubhamsrg):

easiest of all methods would be 1) find probability of drawing out a white marble 2)subtract that from 1 . 3) multiply by 100 to convert to percentage

OpenStudy (anonymous):

probability of taking of white is 12/24=1/2 so probability that a marble drawn at random is NOT white= 1- 1/2=1/2

OpenStudy (anonymous):